# When Benchmarks are Targets: Revealing the Sensitivity of Large Language Model Leaderboards

### v1

@@ 5.5 LLMs readily reference knowledge provided in-context (even if it is misleading)
In our study of in-context knowledge injection we find that LLMs can, expectedly, read off answers to questions when the answer is provided in the context (Table A.27). However when the question is answered incorrectly in the LLM’s context (Table A.26), all models (regardless of size) are unable to reason correctly. This behavior is studied in Wang et al. (2023); Xie et al. (2023) and indicates answer leakage in this way could affect benchmark results.
To test whether LLMs can infer subtler patterns in the few shot examples, we fix all answers in the few-shot examples to each of the positions A/B/C/D. The results (Table 3) suggest that LLMs also bias their answers to these kinds of (potentially inadvertant) patterns in the context.
@@ 6 Related Work
However issues with the leaderboards and the underlying benchmarks have emerged. In a case study, Deng et al. (2023) discovered contamination/leakage of the MMLU benchmark in the training sets of multiple models. A significant portion of models memorized benchmark questions and were able to perfectly reconstruct the removed part of some benchmark questions or asnwers. For instance, GPT-4 correctly completed the questions in 29% of the prompts with URL hinting.
Even under the assumption of uncontaminated data, the performance of models on the underlying benchmarks are not robust to minor perturbations. Pezeshkpour and Hruschka (2023) showed that specific orderings of MMLU answer choices resulted in up to deviations in GPT-4 performance on various subjects. Similarly, Zheng et al. (2023) demonstrate that models are biased to certain answer letters. On llama-30B, they showed a 27% difference in MMLU accuracy by forcing all correct answers to either position A or D. As well, Robinson et al. (2023) find that the accuracy of LLMs improve (without regards to bias) when evaluating using a pure multiple choice question style vs a cloze question answering style.
@@ 7 Conclusion
Building robust leaderboards is a major challenge for the community, as leaderboards help practitioners select the best methods and models for continued research. Given this importance, it is critical to address the break down of existing leaderboards to the slight perturbations we demonstrated in our work. In addition to building our understanding of the causes of this sensitivity (e.g. bias in LLMs and bias in scoring methods), future work should be aimed at adopting and designing benchmark practices that avoid these pitfalls.
@@ 8 Limitations
9 Acknowledgements
We thank Ehsan Hoque for valuable feedback on early versions of the manuscript.
@@ Appendix A.1 Appendix
We present a comprehensive collection of tables containing all the results of our experiments. The often complex nature of the observed behavior warrants a closer look and may inspire novel interpretations for future studies. We believe providing these detailed results will help researchers conduct further analysis and generate hypotheses to help drive research in LLM-benchmarking robustness forward.
@@ A.1.1 Baselines
Model | Acc 0shot | Acc 5shot |
phi-2 | 54.47 | 56.77 |
Yi-6B | 61.12 | 63.23 |
Mistral-7B | 59.56 | 62.36 |
Mistral-7B-Instruct | 53.48 | 53.95 |
Llama-2-7b | 41.81 | 45.88 |
Llama-2-7b-chat | 46.37 | 47.22 |
Llama-2-13b | 52.08 | 55.06 |
Llama-2-13b-chat | 53.12 | 53.53 |
Yi-34B | 73.38 | 76.39 |
Llama-2-70b | 65.44 | 68.78 |
Llama-2-70b-chat | 61.11 | 63.17 |
Model | Acc 0shot | Acc 5shot |
phi-2 | 54.096 | 58.874 |
Yi-6B | 50.512 | 55.034 |
Mistral-7B-v0.1 | 53.584 | 59.556 |
Mistral-7B-Instruct-v0.1 | 52.048 | 54.778 |
Llama-2-7b-hf | 46.331 | 53.072 |
Llama-2-7b-chat-hf | 44.283 | 51.877 |
Llama-2-13b-hf | 48.976 | 56.997 |
Llama-2-13b-chat-hf | 50.256 | 57.594 |
Yi-34B | 61.519 | 64.505 |
Llama-2-70b-hf | 57.253 | 66.126 |
Llama-2-70b-chat-hf | 54.266 | 64.078 |
@@ A.1.2 Answer choice format and ordering
The following tables provide details on the choice formatting manipulation on the selected MMLU subjects.
@@
Yi-6B | 30.66 (-2.00) | 26.68 (3.80) |
Mistral-7B | 34.00 (-1.30) | 22.37 (3.60) |
Mistral-7B-Instruct | 30.33 (-3.70) | 23.08 (-3.90) |
Llama-2-7b | 25.66 (-4.00) | 24.98 (-0.90) |
Llama-2-7b-chat | 28.00 (-3.30) | 28.49 (1.40) |
Llama-2-13b | 29.33 (-5.00) | 26.75 (0.60) |
Llama-2-13b-chat | 29.66 (-4.30) | 20.58 (-1.40) |
Yi-34B | 36.00 (-6.70) | 19.48 (-3.30) |
Llama-2-70b | 37.33 (-2.30) | 10.41 (-4.70) |
Llama-2-70b-chat | 31.66 (-4.30) | 23.25 (-6.30) |
@@ A.1.3 Prompt and scoring modifications
The following tables provide insights on the effect of different scoring styles of MCQs task on MMLU and ARC-C.
@@
Appendix A.2 In-context Knowledge Manipulation
@@
One-shot | Five-shot | |||
Model | Task acc | Acc | Task acc | Acc |
phi-2 | 33.59 | -20.88 | 13.91 | -42.86 |
Yi-6B | 36.13 | -24.99 | 17.97 | -45.26 |
Mistral-7B | 19.51 | -40.05 | 13.20 | -49.16 |
Mistral-7B-Instruct | 10.71 | -42.77 | 4.59 | -49.36 |
Llama-2-7b | 24.25 | -17.56 | 23.63 | -22.25 |
Llama-2-7b-chat | 16.24 | -30.12 | 28.11 | -19.11 |
Llama-2-13b | 12.76 | -39.32 | 4.50 | -50.56 |
Llama-2-13b-chat | 31.49 | -21.63 | 26.30 | -27.22 |
Yi-34B | 32.08 | -41.30 | 37.42 | -38.96 |
Llama-2-70b | 26.27 | -39.17 | 21.54 | -47.24 |
Llama-2-70b-chat | 26.26 | -34.85 | 37.23 | -25.94 |
@@
One-shot | Five-shot | |||
Model | Task Acc | Acc | Task Acc | Acc |
phi-2 | 71.778 | 16.579 | 92.366 | 35.593 |
Yi-6B | 90.91 | 29.23 | 97.09 | 33.86 |
Mistral-7B | 97.45 | 36.85 | 98.99 | 36.63 |
Mistral-7B-Instruct | 98.64 | 45.61 | 99.25 | 45.29 |
Llama-2-7b | 61.00 | 17.68 | 63.82 | 17.94 |
Llama-2-7b-chat | 87.77 | 41.65 | 80.15 | 32.93 |
Llama-2-13b | 96.60 | 43.86 | 99.79 | 44.72 |
Llama-2-13b-chat | 87.02 | 35.11 | 92.69 | 39.17 |
Yi-34B | 99.10 | 23.87 | 98.50 | 22.12 |
Llama-2-70b | 93.45 | 25.75 | 99.09 | 30.31 |
Llama-2-70b-chat | 98.25 | 36.43 | 93.86 | 30.69 |

### v2

@@ 5.5 LLMs readily reference knowledge provided in-context (even if it is misleading)
In our study of in-context knowledge injection, we find that LLMs can, expectedly, read off answers to questions when the answer is provided in the context (Table A.27). However, when the question is answered incorrectly in the LLM’s context (Table A.26), all models (regardless of size) are unable to reason correctly.
To investigate whether this behavior is due to in-context knowledge acquisition or "incorrect answer" pattern following, previous work has shown that smaller models tend to rely more on priors learned during the pretraining stage while larger models tend to be more influenced by knowledge given in-context Wei et al. (2023). This result suggests that we would find that the magnitude of changes in accuracies for small models is smaller than that of larger models. However, we do not observe a significant effect to support that finding conclusively (when comparing A.1 with A.26).
This behavior is further studied in Wang et al. (2023); Xie et al. (2023); Min et al. (2022); Yoo et al. (2022) and indicates answer leakage in this way could affect benchmark results.
To test whether LLMs can infer subtler patterns in the few-shots examples, we fix all answers in the few-shot examples to each of the positions A/B/C/D. The results (Table 3) suggest that LLMs also bias their answers to these kinds of (potentially inadvertent) patterns in the context.
@@ 6 Related Work
However, issues with the leaderboards and the underlying benchmarks have emerged. In a case study, Deng et al. (2023) discovered contamination/leakage of the MMLU benchmark in the training sets of multiple models. A significant portion of models memorized benchmark questions and was able to perfectly reconstruct the removed part of some benchmark questions or answers. For instance, GPT-4 correctly completed the questions in 29% of the prompts with URL hinting.
Even under the assumption of uncontaminated data, the performance of models on the underlying benchmarks are not robust to minor perturbations. Pezeshkpour and Hruschka (2023) showed that specific orderings of MMLU answer choices resulted in up to deviations in GPT-4 performance on various subjects. Similarly, Zheng et al. (2023) demonstrate that models are biased to certain answer letters. On llama-30B, they showed a 27% difference in MMLU accuracy by forcing all correct answers to either position A or D. As well, Robinson et al. (2023) find that the accuracy of LLMs improve (without regard to bias) when evaluating using a pure multiple choice question style vs a cloze question answering style.
@@ 7 Conclusion
Building robust leaderboards is a major challenge for the community, as leaderboards help practitioners select the best methods and models for continued research. Given this importance, it is critical to address the breakdown of existing leaderboards to the slight perturbations we demonstrated in our work. In addition to building our understanding of the causes of this sensitivity (e.g., bias in LLMs and bias in scoring methods), future work should aim to adopt and design benchmark practices that avoid these pitfalls.
@@ 8 Limitations
9 Potential Risks
In this work, we do not present a new leaderboard. There is a risk that Figure 1 is interpreted as a leaderboard or used for model selection. Our intention was to demonstrate how minor changes can affect model rankings.
@@ Appendix A.1 Appendix
We present a comprehensive collection of tables containing the results of all our experiments. The often complex nature of the observed behavior warrants a closer look that may inspire novel interpretations for future studies. We believe providing these detailed results will help researchers conduct further analysis and generate hypotheses to help drive research in LLM-benchmarking robustness forward.
@@ A.1.1 Baselines
Model | Acc 0shot | RStd 0shot | Acc 5shot | RStd 5shot |
phi-2 | 54.47 | 4.01 | 56.77 | 2.65 |
Yi-6B | 61.12 | 3.57 | 63.23 | 2.54 |
Mistral-7B | 59.56 | 4.13 | 62.36 | 1.64 |
Mistral-7B-Instruct | 53.48 | 4.58 | 53.95 | 4.78 |
Llama-2-7b | 41.81 | 8.49 | 45.88 | 8.92 |
Llama-2-7b-chat | 46.37 | 16.11 | 47.22 | 12.15 |
Llama-2-13b | 52.08 | 12.04 | 55.06 | 4.42 |
Llama-2-13b-chat | 53.12 | 12.80 | 53.53 | 8.32 |
Yi-34B | 73.38 | 5.17 | 76.39 | 2.16 |
Llama-2-70b | 65.44 | 3.20 | 68.78 | 1.56 |
Llama-2-70b-chat | 61.11 | 10.95 | 63.17 | 8.06 |
Model | Acc 0shot | RStd 0shot | Acc 5shot | RStd 5shot |
phi-2 | 54.096 | 2.558 | 58.874 | 2.509 |
Yi-6B | 50.512 | 2.114 | 55.034 | 0.737 |
Mistral-7B | 53.584 | 2.578 | 59.556 | 1.037 |
Mistral-7B-Instruct | 52.048 | 1.443 | 54.778 | 2.022 |
Llama-2-7b | 46.331 | 4.094 | 53.072 | 0.837 |
Llama-2-7b-hf | 44.283 | 2.175 | 51.877 | 1.399 |
Llama-2-13b | 48.976 | 2.923 | 56.997 | 0.799 |
Llama-2-13b-chat | 50.256 | 1.841 | 57.594 | 2.991 |
Yi-34B | 61.519 | 2.537 | 64.505 | 1.48 |
Llama-2-70b | 57.253 | 2.657 | 66.126 | 1.926 |
Llama-2-70b-chat | 54.266 | 1.505 | 64.078 | 2.084 |
@@ A.1.2 Answer choice format and ordering
The following tables provide details on the choice formatting manipulation on the three selected MMLU subjects.
@@
Yi-6B | 30.66 (-2.0) | 26.68 (3.8) |
Mistral-7B | 34.00 (-1.3) | 22.37 (3.6) |
Mistral-7B-Instruct | 30.33 (-3.7) | 23.08 (-3.9) |
Llama-2-7b | 25.66 (-4.0) | 24.98 (-0.9) |
Llama-2-7b-chat | 28.00 (-3.3) | 28.49 (1.4) |
Llama-2-13b | 29.33 (-5.0) | 26.75 (0.6) |
Llama-2-13b-chat | 29.66 (-4.3) | 20.58 (-1.4) |
Yi-34B | 36.00 (-6.7) | 19.48 (-3.3) |
Llama-2-70b | 37.33 (-2.3) | 10.41 (-4.7) |
Llama-2-70b-chat | 31.66 (-4.3) | 23.25 (-6.3) |
@@ A.1.3 Prompt and scoring modifications
The following tables provide results on the effect of different scoring styles of MCQs task on MMLU and ARC-C.
@@
A.1.4 In-context Knowledge Manipulation
@@
Model | Task Acc | Task Acc |
1-shot | 5-shot | |
phi-2 | 33.59 | 13.91 |
Yi-6B | 36.13 | 17.97 |
Mistral-7B | 19.51 | 13.20 |
Mistral-7B-Instruct | 10.71 | 4.59 |
Llama-2-7b | 24.25 | 23.63 |
Llama-2-7b-chat | 16.24 | 28.11 |
Llama-2-13b | 12.76 | 4.50 |
Llama-2-13b-chat | 31.49 | 26.30 |
Yi-34B | 32.08 | 37.42 |
Llama-2-70b | 26.27 | 21.54 |
Llama-2-70b-chat | 26.26 | 37.23 |
@@
Task Acc | Task Acc | |
Model | 1-shot | 5-shot |
phi-2 | 71.778 | 92.366 |
Yi-6B | 90.91 | 97.09 |
Mistral-7B | 97.45 | 98.99 |
Mistral-7B-Instruct | 98.64 | 99.25 |
Llama-2-7b | 61.00 | 63.82 |
Llama-2-7b-chat | 87.77 | 80.15 |
Llama-2-13b | 96.60 | 99.79 |
Llama-2-13b-chat | 87.02 | 92.69 |
Yi-34B | 99.10 | 98.50 |
Llama-2-70b | 93.45 | 99.09 |
Llama-2-70b-chat | 98.25 | 93.86 |
@@
A.1.5 MMLU Overview
subject | choice_A | choice_B | choice_C | choice_D | total |
abstract_algebra | 22 | 26 | 31 | 21 | 100 |
anatomy | 25 | 34 | 45 | 31 | 135 |
astronomy | 27 | 28 | 46 | 51 | 152 |
business_ethics | 30 | 26 | 23 | 21 | 100 |
clinical_knowledge | 57 | 71 | 58 | 79 | 265 |
college_biology | 37 | 32 | 37 | 38 | 144 |
college_chemistry | 20 | 21 | 18 | 41 | 100 |
college_computer_science | 26 | 15 | 26 | 33 | 100 |
college_mathematics | 21 | 23 | 25 | 31 | 100 |
college_medicine | 36 | 36 | 43 | 58 | 173 |
college_physics | 22 | 20 | 22 | 38 | 102 |
computer_security | 28 | 24 | 30 | 18 | 100 |
conceptual_physics | 62 | 76 | 48 | 49 | 235 |
econometrics | 27 | 32 | 28 | 27 | 114 |
electrical_engineering | 35 | 32 | 43 | 35 | 145 |
elementary_mathematics | 79 | 97 | 101 | 101 | 378 |
formal_logic | 36 | 25 | 19 | 46 | 126 |
global_facts | 18 | 31 | 33 | 18 | 100 |
high_school_biology | 55 | 79 | 78 | 98 | 310 |
high_school_chemistry | 31 | 55 | 60 | 57 | 203 |
high_school_computer_science | 25 | 23 | 33 | 19 | 100 |
high_school_european_history | 36 | 40 | 47 | 42 | 165 |
high_school_geography | 35 | 43 | 50 | 70 | 198 |
high_school_government_and_politics | 38 | 40 | 44 | 71 | 193 |
high_school_macroeconomics | 79 | 86 | 83 | 142 | 390 |
high_school_mathematics | 57 | 71 | 71 | 71 | 270 |
high_school_microeconomics | 50 | 55 | 50 | 83 | 238 |
high_school_physics | 30 | 30 | 41 | 50 | 151 |
high_school_psychology | 105 | 129 | 121 | 190 | 545 |
high_school_statistics | 33 | 35 | 46 | 102 | 216 |
high_school_us_history | 51 | 48 | 53 | 52 | 204 |
high_school_world_history | 64 | 62 | 63 | 48 | 237 |
human_aging | 70 | 84 | 45 | 24 | 223 |
human_sexuality | 34 | 30 | 30 | 37 | 131 |
international_law | 29 | 30 | 45 | 17 | 121 |
jurisprudence | 28 | 32 | 25 | 23 | 108 |
logical_fallacies | 36 | 40 | 49 | 38 | 163 |
machine_learning | 35 | 32 | 27 | 18 | 112 |
management | 18 | 26 | 20 | 39 | 103 |
marketing | 68 | 60 | 60 | 46 | 234 |
medical_genetics | 30 | 26 | 20 | 24 | 100 |
miscellaneous | 186 | 225 | 212 | 160 | 783 |
moral_disputes | 86 | 85 | 101 | 74 | 346 |
moral_scenarios | 213 | 217 | 221 | 244 | 895 |
nutrition | 69 | 70 | 77 | 90 | 306 |
philosophy | 58 | 85 | 93 | 75 | 311 |
prehistory | 70 | 86 | 95 | 73 | 324 |
professional_accounting | 66 | 72 | 76 | 68 | 282 |
professional_law | 377 | 367 | 415 | 375 | 1534 |
professional_medicine | 50 | 55 | 45 | 122 | 272 |
professional_psychology | 153 | 157 | 169 | 133 | 612 |
public_relations | 24 | 38 | 23 | 25 | 110 |
security_studies | 46 | 42 | 59 | 98 | 245 |
sociology | 49 | 48 | 50 | 54 | 201 |
us_foreign_policy | 28 | 21 | 25 | 26 | 100 |
virology | 47 | 53 | 34 | 32 | 166 |
world_religions | 55 | 36 | 50 | 30 | 171 |
total | 3222 | 3462 | 3582 | 3776 | 14042 |
subject | choice_A | choice_B | choice_C | choice_D | total |
abstract_algebra | 2 | 2 | 1 | 0 | 5 |
anatomy | 0 | 1 | 2 | 2 | 5 |
astronomy | 3 | 0 | 0 | 2 | 5 |
business_ethics | 1 | 1 | 1 | 2 | 5 |
clinical_knowledge | 1 | 2 | 2 | 0 | 5 |
college_biology | 0 | 3 | 1 | 1 | 5 |
college_chemistry | 2 | 0 | 1 | 2 | 5 |
college_computer_science | 0 | 2 | 0 | 3 | 5 |
college_mathematics | 0 | 2 | 1 | 2 | 5 |
college_medicine | 2 | 1 | 1 | 1 | 5 |
college_physics | 3 | 1 | 0 | 1 | 5 |
computer_security | 1 | 1 | 1 | 2 | 5 |
conceptual_physics | 2 | 1 | 2 | 0 | 5 |
econometrics | 1 | 0 | 3 | 1 | 5 |
electrical_engineering | 1 | 2 | 1 | 1 | 5 |
elementary_mathematics | 1 | 4 | 0 | 0 | 5 |
formal_logic | 0 | 1 | 3 | 1 | 5 |
global_facts | 2 | 3 | 0 | 0 | 5 |
high_school_biology | 1 | 0 | 1 | 3 | 5 |
high_school_chemistry | 1 | 0 | 3 | 1 | 5 |
high_school_computer_science | 0 | 1 | 3 | 1 | 5 |
high_school_european_history | 2 | 1 | 1 | 1 | 5 |
high_school_geography | 1 | 2 | 1 | 1 | 5 |
high_school_government_and_politics | 1 | 0 | 2 | 2 | 5 |
high_school_macroeconomics | 0 | 0 | 3 | 2 | 5 |
high_school_mathematics | 0 | 1 | 2 | 2 | 5 |
high_school_microeconomics | 0 | 1 | 2 | 2 | 5 |
high_school_physics | 0 | 2 | 0 | 3 | 5 |
high_school_psychology | 1 | 2 | 1 | 1 | 5 |
high_school_statistics | 0 | 0 | 1 | 4 | 5 |
high_school_us_history | 0 | 2 | 2 | 1 | 5 |
high_school_world_history | 1 | 3 | 0 | 1 | 5 |
human_aging | 1 | 2 | 2 | 0 | 5 |
human_sexuality | 3 | 1 | 1 | 0 | 5 |
international_law | 2 | 2 | 1 | 0 | 5 |
jurisprudence | 2 | 0 | 0 | 3 | 5 |
logical_fallacies | 0 | 1 | 2 | 2 | 5 |
machine_learning | 1 | 1 | 2 | 1 | 5 |
management | 3 | 0 | 1 | 1 | 5 |
marketing | 1 | 1 | 0 | 3 | 5 |
medical_genetics | 3 | 0 | 1 | 1 | 5 |
miscellaneous | 1 | 2 | 2 | 0 | 5 |
moral_disputes | 3 | 1 | 0 | 1 | 5 |
moral_scenarios | 1 | 1 | 2 | 1 | 5 |
nutrition | 1 | 1 | 2 | 1 | 5 |
philosophy | 1 | 0 | 3 | 1 | 5 |
prehistory | 2 | 1 | 1 | 1 | 5 |
professional_accounting | 2 | 2 | 0 | 1 | 5 |
professional_law | 2 | 2 | 0 | 1 | 5 |
professional_medicine | 0 | 0 | 1 | 4 | 5 |
professional_psychology | 2 | 0 | 0 | 3 | 5 |
public_relations | 1 | 0 | 2 | 2 | 5 |
security_studies | 0 | 2 | 2 | 1 | 5 |
sociology | 1 | 3 | 1 | 0 | 5 |
us_foreign_policy | 1 | 0 | 2 | 2 | 5 |
virology | 2 | 1 | 1 | 1 | 5 |
world_religions | 1 | 3 | 0 | 1 | 5 |
total | 67 | 69 | 71 | 78 | 285 |